Factor completely:
x^4 - 6x^2 y^2 + 9y^4
1
Expert's answer
2013-02-20T11:29:23-0500
Let's use the perfect square formula:

x^4 - 6x²y² + 9y^4 = (x² - 3y²)2

And now let's use the difference of two squares formula:

(x² - 3y²) = (x - √3y)(x + √3y).

Now factorization is complete. Therefore,

x^4 - 6x^2 y^2 + 9y^4 = (x - √3y)2(x + √3y)2